OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (2, 4) and a directrix of y = 8

OpenStudy (gnorris):

multiple choice?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

@gnorris

OpenStudy (anonymous):

f(x) = −one eighth (x − 2)^2 + 6 f(x) = one eighth (x + 2)^2 + 8 f(x) = one eighth (x − 2)^2 + 6 f(x) = −one eighth (x + 2^)2 + 8

OpenStudy (gnorris):

alright let me see if i can

OpenStudy (gnorris):

f(x) = -1/8 (x - 2)2 + 6 f(x) = 1/8 (x - 2)2 + 6 f(x) = -1/8 (x + 2)2 + 8 f(x) = 1/8 (x + 2)2 + 8

OpenStudy (gnorris):

i wrote it out better

OpenStudy (anonymous):

yeah

OpenStudy (gnorris):

urghh i cant be much help all i can say is its not the third choice

OpenStudy (anonymous):

really that is the one i thought it was

OpenStudy (anonymous):

The \[\frac{ 1 }{ 4p }\] is the most confusing part for me

OpenStudy (gnorris):

no it isn't the third choice Go through the steps. Focus is at (2,4), directrix is at y = 8. Distance between focus and directrix is |8-4| = 4. Vertex (h,k) is at spot halfway between focus and directrix, or (h,k)=(2,8+4/2)=(2,6) Fill in the blanks: y=a(x−h)2+k y=a(x−2)2+6 To find the value of a, use a=1/4p where p is the distance from the focus (2,4) to the vertex (2,6). draw a sketch; the parabola wraps around the focus and veers away from the directrix (and does not cross it!) If the parabola opens upward, like a bowl, a will be greater than 0. If the parabola opens downward, like an inverted bowl, a will be less than 0.

OpenStudy (anonymous):

uhh ok that makes a little more sense

OpenStudy (anonymous):

could you use the distance formula

OpenStudy (anonymous):

to find p i mean